OpenStudy (superfly123):

Find a value of theta for which the statement is true. Theta will represent x here. tan x = cot5x

OpenStudy (anonymous):

the above equation can become cot5x * cotx = 1 or cos5x*cosx - sin5x * sinx =0 Does this help ?

OpenStudy (superfly123):

Idk all i did was x plus 5x equals 90

OpenStudy (superfly123):

got x equals 15 dont know hat to do next

Directrix (directrix):

Is this: cot5x cot^5 x or is it cotangent of (5*x) ?

OpenStudy (superfly123):

ill show you on my paper

Directrix (directrix):

Okay

OpenStudy (anonymous):

not just 90 but \[\pm90+360k\] where k is any integer

OpenStudy (anonymous):

or in pi notation: \[\pm \frac{ \pi }{ 2 } + 2\pi k\]

OpenStudy (superfly123):

Its number 18

OpenStudy (anonymous):

if it is just one value then \[\pi /12\] or 15 degrees is fine

Directrix (directrix):

Thanks for posting that. @SuperFly123

OpenStudy (superfly123):

xavier i dont understand what your doing sorry

OpenStudy (superfly123):

No prob @Directrix

OpenStudy (anonymous):

@SuperFly123 - i thought you needed a generalised solution not just one angle :) Anyhow, for one acute angle your answer should be correct

OpenStudy (phi):

I think you can use an addition formula to simplify the problem. first change things in sin and cos

OpenStudy (superfly123):

But i think i need to find the exact value

OpenStudy (superfly123):

phi, how?

Directrix (directrix):

@SuperFly123 Where in the first quadrant is tan(θ) = 1 ?

OpenStudy (phi):

write tan and cot in terms of sin and cos. do you know how to do that ?

OpenStudy (anonymous):

@SuperFly123 - Can you show your approach ?

OpenStudy (superfly123):

no phi, and ill show you how my teacher did an example

OpenStudy (phi):

ok, do that, because there are probably a few different ways to tackle this.

OpenStudy (pawanyadav):

tan x= cot 5x= tan (90°-5x) So x=90-5x 6x=90 x=15° ,,

OpenStudy (phi):

yes, that's pretty.

OpenStudy (phi):

\[ \frac{\sin x}{\cos x}= \frac{\cos 5x }{\sin 5x} \\ \sin x \sin 5x = \cos 5x \cos x \\ \cos 5x \cos x - \sin5x \sin x =0\\ \cos(6x)= 0\]

OpenStudy (pawanyadav):

I think , it can have more than one solutions.

Directrix (directrix):

>>Find a value of theta The question sounds as if one value will suffice.

OpenStudy (phi):

yes, my approach might be overkill

OpenStudy (pawanyadav):

Sure

OpenStudy (superfly123):

I got 15 what do i do from there

Directrix (directrix):

How about a basic approach of knowing that tan 45 = 1 and knowing that 5*45 = 225 degrees, the cotangent of which is 1. That gives π/4 as an angle that satisfies the equation.

OpenStudy (superfly123):

But, where do we get 45 from

Directrix (directrix):

From the basic trig values I had to learn.

Directrix (directrix):

The table comes from right triangle trig.

OpenStudy (superfly123):

Since, my teacher didnt get up to that yet i dont think i should use that approach

Directrix (directrix):

Okay.

OpenStudy (superfly123):

I dont know what to do omg

OpenStudy (superfly123):

What i have so far is tan 15 = cot 75 Then tan 15 = 1/tan(45+30) giving me a right answer

OpenStudy (phi):

you should fix question 4. The "trick" is if you are told a function and it's "cofunction" are equal then the sum of the two angles is 90 degrees for example, if you are told sin (A) = cos(B) then A+B= 90 or if tan A = cot B then A+B= 90

OpenStudy (superfly123):

Yeh i know i got that wrong

OpenStudy (phi):

so in question 4, you don't set the two expressions equal to each other. you add them up, and set that sum equal to 90

OpenStudy (phi):

for the tan x = cot (5x) you do x+5x= 90 6x= 90 x= 15 which gives one (of multiple possible) answers

OpenStudy (superfly123):

yeh so the answer is just 15

OpenStudy (anonymous):

yes :)

OpenStudy (superfly123):

Alright then

OpenStudy (phi):

can you do these problems? for example 13: sin 10 = cos x they are "co functions" so you can write 10+x= 90 x= 80 (using a calculator we can check sin 10 = cos 80)